# Unified Multi-Rate Model Predictive Control for a Jet-Powered Humanoid Robot

### v1

@@ II-A Notation
  • ${}^{\mathcal{A}}R_{\mathcal{B}}$ is the rotation matrix from frame $\mathcal{B}$ to frame $\mathcal{B}$;
@@ II-B Floating-base model
A flying humanoid robot can be modelled as a multi-body system composed of $n+1$ rigid bodies, called links, connected by $n$ joints with one degree of freedom. Following the floating base formalism, it is possible to define the configuration space $\mathbb{Q}\in\mathbb{R}^{3}\times SO(3)\times\mathbb{R}^{n}$; an element of the space is composed of the tuple $q=({}^{\mathcal{J}}p_{\mathcal{B}},{}^{\mathcal{J}}R_{\mathcal{B}},s)$. The velocity belongs to the space $\mathbb{V}\in\mathbb{R}^{3}\times\mathbb{R}^{3}\times\mathbb{R}^{n}$ and an element of the space $\mathbb{V}$ is $\nu=({}^{\mathcal{J}}\dot{p}_{\mathcal{B}},{}^{\mathcal{J}}\omega_{\mathcal{B}},\dot{s})$. By applying the Euler-Poincaré formalism, the equation of motion for a multi-body system subject to m external forces becomes [27]:
@@ III-A Centroidal momentum equations
${{}^{\mathcal{B}}R_{\mathcal{J}}}{{}^{\mathcal{\mathcal{G}[\mathcal{J}]}}\dot{h}^{p}}=$ $A_{ang}(s)T,$ (6)
where ${T}$ is the vector collecting the thrust forces ${T=[T_{1},\dots,T_{n_{j}}]}$ of the $n_{j}$ jet turbines, the matrix ${A}$ is defined as:
@@
under the assumption that ${{}^{\mathcal{B}}\omega_{\mathcal{G}}}\approx{{}^{\mathcal{B}}\omega_{\mathcal{B}}}$. This approximation holds when the joint velocities are sufficiently small, a condition that is satisfied during flight, as the robot moves its arms slowly to adjust the jet directions [17]. With this assumption, the relationship between the derivative of the Euler angles $\dot{\phi}$ and the angular momentum ${}^{\mathcal{\mathcal{G}[\mathcal{B}]}}h^{w}$ can be expressed as:
@@ III-B Dynamical model
$\begin{cases}{}^{\mathcal{J}}\dot{x}_{\mathcal{G}}=\frac{1}{m}{{}^{\mathcal{J}}R_{\mathcal{B}}}{{}^{\mathcal{\mathcal{G}[\mathcal{B}]}}h^{p}}\\[3.0pt] {}^{\mathcal{\mathcal{G}[\mathcal{B}]}}\dot{h}^{p}=A_{lin}(s)T+mg{{}^{\mathcal{B}}R_{\mathcal{J}}}e_{3}-S(^{\mathcal{B}}\omega_{\mathcal{B}}){{{}^{\mathcal{\mathcal{G}[\mathcal{B}]}}h^{p}}}\\[3.0pt] \dot{\phi}={({{}^{\mathcal{\mathcal{G}[\mathcal{B}]}}\mathbb{I}}W)^{-1}}{{}^{\mathcal{B}}R_{\mathcal{J}}}{{}^{\mathcal{\mathcal{G}[\mathcal{B}]}}h^{w}}\\[3.0pt] {}^{\mathcal{\mathcal{G}[\mathcal{B}]}}\dot{h}^{w}={A}_{ang}({s}){T}-S(^{\mathcal{B}}\omega_{\mathcal{B}}){{}^{\mathcal{\mathcal{G}[\mathcal{B}]}}h^{w}}\\[3.0pt] \dot{T}_{i}=\dot{T}_{i}\\[3.0pt] \ddot{T}_{i}=h(T_{i},\dot{T}_{i})+g(T_{i},\dot{T}_{i}){v_{i}}\\[3.0pt] \end{cases}$ (15)
@@ III-C Linearised robot dynamics
$f(z,u)\approx f(z_{c},u_{c})+\left.\frac{\partial f}{\partial z}\right|_{z_{c},u_{c}}(z-z_{c})+\left.\frac{\partial f}{\partial u}\right|_{z_{c},u_{c}}(u-u_{c}).$
@@
$\begin{cases}\dot{x}_{\mathcal{G}}\approx\frac{1}{m}{{}^{\mathcal{J}}R_{\mathcal{B}}}l\\[3.0pt] \dot{h}^{p}\approx{A}_{lin}(s_{c})T+mg{{}^{\mathcal{B}}R_{\mathcal{J}}}e_{3}+\lambda_{lin}(s-s_{c})-S(\omega_{\mathcal{B}})h^{p}\\[3.0pt] \dot{\phi}\approx\left({\mathbb{I}}W\right)^{-1}h^{w}\\[3.0pt] \dot{h}^{w}\approx{A}_{ang}(s_{c})T+\lambda_{ang}(s-s_{c})-S(\omega_{\mathcal{B}})h^{w}\\[3.0pt] \dot{T}_{i}=\dot{T}_{i}\\[3.0pt] \begin{aligned} \ddot{T}_{i}\approx h(T_{c,i},\dot{T}_{c,i})+\frac{\partial(h+gv)}{\partial T}\bigg{|}_{T_{c,i},\dot{T}_{c,i},v_{c,i}}(T_{i}-T_{c,i})\\ +\frac{\partial(h+gv)}{\partial\dot{T}}\bigg{|}_{T_{c,i},\dot{T}_{c,i},v_{c,i}}(\dot{T}_{i}-\dot{T}_{c,i})+g(T_{c,i},\dot{T}_{c,i})v_{i}\end{aligned}\end{cases}$ (17)
@@
where $c(q,u)$ represents the bias terms due to the linearisation. For the sake of clarity, in the following chapters we will refer to the matrices $A(z,u)$ and $B(z,u)$ and to the vector $c(z,u)$ evaluated in the current state of the robot as $A$, $B$ and $c$.
@@ III-D MPC formulation
In order to improve the tracking of a reference in position and orientation, the state $z$ is augmented to include two additional error variables $e_{x}$ and $e_{\phi}$ [29]; the effect of adding those variables is to penalise the steady state offset. The augmented state thus becomes:
@@ III-E Multi-rate MPC
Let us assume that the MPC runs at a frequency $f_{\text{MPC}}$ and that a given actuator operates at a lower frequency $f_{s}$, with $f_{\text{MPC}}>f_{s}$. Furthermore, suppose that $f_{\text{MPC}}$ is an integer multiple of $f_{s}$, such that $N_{r}=\frac{f_{\text{MPC}}}{f_{s}}$. In this case, the number of control inputs $N_{u_{s}}$ associated with the actuator within the prediction horizon is chosen as $N_{u_{s}}=\frac{N_{h}}{N_{r}}$, ensuring that the control input timing matches the actuator’s update frequency. Figure 2 shows the variation of two different control inputs with two different control frequencies along the prediction horizon of the MPC. The red line represents a control input $u_{f}$ that is updated at every propagation step of the MPC, while the blue line represents a control input $u_{s}$ that is updated every $N_{r}$ steps of the MPC.
[…]
@@
0:  $f_{\text{MPC}}$, $f_{s}$, $N_{r}$, $u_{\text{prev}},j$
1:  $N_{r}\leftarrow f_{\text{MPC}}/f_{s}$
2:  $j\leftarrow N_{r}$
@@ IV-A Simulation Environment
The simulation environment used to perform the experiments is MuJoCo [30], running at a frequency of 1000 Hz. All simulations are conducted on an Ubuntu 22.04 operating system equipped with an Intel(R) Core(TM) i7-12700H processor.
[…]
@@
In our case, the low-level controller of the joints operates at 1000 Hz, while the low-level controller of the jet engines operates at 10 Hz. Those values are chosen taking into consideration the specifics of the robot iRonCub. To enable the MPC to run at a higher frequency, we adopt the approach proposed in [31], which adapts the sampling intervals of the MPC along the prediction horizon, with a shorter sampling time at the beginning of the horizon, which progressively increases towards the end. In fact, given the slow dynamics of the jet engines, we chose a control horizon of 1 second. Given that control horizon, to make the MPC run at 200 Hz with a constant timestep, we would need 200 knots, which would considerably slow down the problem. Exploiting this method, we can run the MPC at a frequency of 200 Hz with 17 knots. Recalling the notation in Subsection III-E, the low-frequency control $u_{s}$ input is the jet turbines control input $v$, which varies at a frequency of 10 Hz. The joint positions $s$ are computed at every step of the MPC, and the number $N_{r}$, since we adopt a variable sampling, varies along the control horizon, but it is chosen such that the jet turbine throttle is kept constant for 0,1 seconds.
@@ IV-B1 Disturbance recovery
[…]
@@ IV-B2 Minimum Jerk Trajectory Tracking
[…]
@@ IV-D1 Jet Dynamics study
IV-D2 Multi-rate study
@@ V-A Linearization
Exploiting the property $S(x)y=-S(y)x$ and writing ${}^{B}\omega_{B,J_{j}}=J^{rel}_{\omega_{j}}(s)\dot{s}$ we can finally rewrite:
@@
where ${}^{B}r_{i}$ is the distance between the centre of mass and the $i^{th}$ turbine expressed in body frame. Exploiting once again the definition of the derivative of the rotation matrix, the property $S(x)y=-S(y)x$ and using the linear and angular Jacobians ${}^{B}\omega_{B,J_{j}}=J^{rel}_{\omega_{j}}(s)\dot{s}$ and ${}^{B}\omega_{B,J_{j}}=J^{rel}_{\omega_{j}}(s)\dot{s}$, (27) rewrites as:

### v2

@@ II-A Notation
  • $^{\mathcal{A}}R_{\mathcal{B}}$ is the rotation matrix from frame $\mathcal{A}$ to frame $\mathcal{B}$;
@@ II-B Floating-base model
A flying humanoid robot can be modelled as a multi-body system composed of $n+1$ rigid bodies, called links, connected by $n$ joints with one degree of freedom. Following the floating base formalism, it is possible to define the configuration space $\mathbb{Q}\in\mathbb{R}^{3}\times SO(3)\times\mathbb{R}^{n}$; an element of the space is composed of the tuple $q=({}^{\mathcal{J}}p_{\mathcal{J},\mathcal{B}},{}^{\mathcal{J}}R_{\mathcal{B}},s)$. The velocity belongs to the space $\mathbb{V}\in\mathbb{R}^{3}\times\mathbb{R}^{3}\times\mathbb{R}^{n}$ and an element of the space $\mathbb{V}$ is $\nu=({}^{\mathcal{J}}\dot{p}_{\mathcal{J},\mathcal{B}},{}^{\mathcal{J}}\omega_{\mathcal{B}},\dot{s})$. By applying the Euler-Poincaré formalism, the equation of motion for a multi-body system subject to m external forces becomes [27]:
@@ III-A Centroidal momentum equations
${{}^{\mathcal{B}}R_{\mathcal{J}}}{{}^{\mathcal{\mathcal{G}[\mathcal{J}]}}\dot{h}^{w}}$= $A_{ang}(s)T,$ (6)
where $s$ represents the joint positions and T{T} is the vector collecting the thrust forces $T=[T_{1},\dots,T_{n_{j}}]$} of the $n_{j}$ jet turbines, the matrix A{A} is defined as:
@@
under the assumption that ${{}^{\mathcal{B}}\omega_{\mathcal{G}}}\approx{{}^{\mathcal{B}}\omega_{\mathcal{B}}}$. This approximation holds when the joint velocities are sufficiently small, a condition that is satisfied during flight, as the robot moves its arms slowly to adjust the jet directions [17]. With this assumption, the relationship between the derivative of the Euler angles $\dot{\phi}\in\mathbb{R}^{3}$ and the angular momentum $^{\mathcal{\mathcal{G}[\mathcal{B}]}}h^{w}$ can be expressed as:
@@ III-B Dynamical model
[…]
@@ III-C Linearised robot dynamics
$f(z,u)\approx f(z_{c},u_{c})+\left.\frac{\partial f}{\partial z}\right|_{z_{c},u_{c}}\hskip-8.99994pt(z-z_{c})+\left.\frac{\partial f}{\partial u}\right|_{z_{c},u_{c}}\hskip-8.99994pt(u-u_{c}).$
@@
$\begin{cases}\dot{x}_{\mathcal{G}}\approx\frac{1}{m}{{}^{\mathcal{J}}R_{\mathcal{B}}}h^{p}\\[3.0pt] \dot{h}^{p}\approx{A}_{lin}(s_{c})T+mg{{}^{\mathcal{B}}R_{\mathcal{J}}}e_{3}+\lambda_{lin}(s-s_{c})-S(\omega_{\mathcal{B}})h^{p}\\[3.0pt] \dot{\phi}\approx\left({\mathbb{I}}W\right)^{-1}h^{w}\\[3.0pt] \dot{h}^{w}\approx{A}_{ang}(s_{c})T+\lambda_{ang}(s-s_{c})-S(\omega_{\mathcal{B}})h^{w}\\[3.0pt] \dot{T}_{i}=\dot{T}_{i}\\[3.0pt] \begin{aligned} \ddot{T}_{i}\approx h(T_{c,i},\dot{T}_{c,i})+\frac{\partial(h+gv)}{\partial T}\bigg{|}_{T_{c,i},\dot{T}_{c,i},v_{c,i}}(T_{i}-T_{c,i})\\ +\frac{\partial(h+gv)}{\partial\dot{T}}\bigg{|}_{T_{c,i},\dot{T}_{c,i},v_{c,i}}(\dot{T}_{i}-\dot{T}_{c,i})+g(T_{c,i},\dot{T}_{c,i})v_{i}\end{aligned}\end{cases}$ (17)
@@
where $c(z,u)$ represents the bias terms due to the linearisation. For the sake of clarity, in the following chapters we will refer to the matrices $A(z,u)$ and $B(z,u)$ and to the vector $c(z,u)$ evaluated in the current state of the robot as $A$, $B$ and $c$.
@@ III-D MPC formulation
In order to improve the tracking of a reference in position and orientation, the state $z$ is augmented to include two additional integral error variables $e_{x}$ and $e_{\phi}$ [29]; the effect of adding those variables is to penalise the steady state offset. The augmented state thus becomes:
@@ III-E Multi-rate MPC
Let us assume that the MPC runs at a frequency $f_{\text{MPC}}$ and that a given actuator operates at a lower frequency $f_{s}$, with $f_{\text{MPC}}>f_{s}$. Furthermore, suppose that $f_{\text{MPC}}$ is an integer multiple of $f_{s}$, such that $N_{r}=\frac{f_{\text{MPC}}}{f_{s}}$. In this case, the number of control inputs $N_{u_{s}}$ associated with the actuator within the prediction horizon is chosen as $N_{u_{s}}=\frac{N_{h}}{N_{r}}$, where $N_{h}$ is the horizon of the MPC, ensuring that the control input timing matches the actuator’s update frequency. Figure 2 shows the variation of two different control inputs with two different control frequencies along the prediction horizon of the MPC. The red line represents a control input $u_{f}$ that is updated at every propagation step of the MPC, while the blue line represents a control input $u_{s}$ that is updated every $N_{r}$ steps of the MPC.
$\color[rgb]{0.29,0.56,0.89}\definecolor[named]{pgfstrokecolor}{rgb}{0.29,0.56,0.89}u_{s}$$\color[rgb]{0.82,0.01,0.11}\definecolor[named]{pgfstrokecolor}{rgb}{0.82,0.01,0.11}u_{f}$$N_{r}$$N_{h}$KnotsKnotsInputsInputs
@@
[…]
@@ IV-A Simulation Environment
The simulation environment used to perform the experiments is MuJoCo [30], running at a frequency of 1000 Hz. All simulations are conducted on an Ubuntu 22.04 operating system equipped with an Intel(R) Core(TM) i7-12700H processor [^1] The code to run the simulations is available at https://github.com/ami-iit/paper_gorbani_2025_humanoids_multi-rate-mpc-ironcub. In order to have a simulation that is closer to reality, the jet dynamics are simulated using the neural network model proposed in [22], which is identified using data gathered from the real turbines. To identify the second-order model (12), the Extended Kalman Filter (EKF) algorithm described in [19] is employed. Figure 3 presents a comparison between the thrust predicted by the neural network model and the thrust generated by the identified non-linear model, when both are provided with the same throttle input profile. The mean absolute error (MAE) between the two models is approximately $5.95~\text{N}$. This comparison highlights the mismatch between the model used in simulation and one used in the prediction model of the MPC.
@@
In our case, the low-level controller of the joints operates at 1000 Hz, while the low-level controller of the jet engines operates at 10 Hz. Those values are chosen taking into consideration the specifics of the robot iRonCub. To enable the MPC to run at a higher frequency, we adopt the approach proposed in [31], which adapts the sampling intervals of the MPC along the prediction horizon, with a shorter sampling time at the beginning of the horizon, which progressively increases towards the end. In fact, given the slow dynamics of the jet engines, we chose a control horizon of 1 second. Given that control horizon, to make the MPC run at 200 Hz with a constant timestep, we would need 200 knots, which would considerably slow down the problem. Exploiting this method, we can run the MPC at a frequency of 200 Hz with 17 knots. Recalling the notation in Subsection III-E, the low-frequency control $u_{s}$ input is the jet turbines control input $v$, which varies at a frequency of 10 Hz. The joint positions $s$ are computed at every step of the MPC, and the number $N_{r}$, since we adopt a variable sampling, varies along the control horizon, but it is chosen such that the jet turbine throttle is kept constant for $0.1$ seconds.
@@ IV-B1 Disturbance recovery
To demonstrate the controller’s ability to reject external disturbances, an external torque of $300~\text{Nm}$ was applied along the $y$-axis, and an external force of $50~\text{N}$ was applied along the $x$-axis of the root link for $0.1~\text{seconds}$ at $t=2~\text{s}$. As shown in Figure 4, the robot successfully recovers from the disturbance, despite experiencing a tilt of approximately $15^{\circ}$ in the roll axis, a displacement of about $1~\text{m}$ along the $x$-axis, and a vertical drop of approximately $1~\text{m}$.
@@ IV-B2 Minimum Jerk Trajectory Tracking
To test the performance of the MPC in tracking a reference trajectory, a minimum jerk trajectory is generated by interpolating a given set of points and used as the reference for the centre of mass position $^{\mathcal{J}}x_{\mathcal{G},ref}$}, while the reference orientation $\phi_{ref}$ is constant. The reference linear momentum is computed as $h^{l}_{ref}=m{{}^{\mathcal{B}}R_{\mathcal{J}}}{{}^{\mathcal{J}}\dot{x}_{\mathcal{G},ref}}$, while the reference angular momentum $h^{w}_{ref}$ is set to zero since the reference orientation is constant. Figure 5 shows the tracking of the centre of mass and the orientation of the base. The blue line is the measured signal, while the green one is the reference; the $x$ and $y$ components of the position are tracked fairly well by the controller, while the $z$ component presents an offset of about $0.15~\text{m}$. Concerning the orientation, on Roll and Yaw have a Mean Absolute Error lower than $0.01$ rad, while for the Pitch it is about $0.03$ rad. The error in the tracking is due to the discrepancies between the neural network model used to simulate the jet dynamics and the linearised non-linear model used in the prediction model of the MPC.
@@ IV-D1 Jet Dynamics study
IV-D2 LPV study
To demonstrate the relevance of the LPV formulation, we carried out a simulation in which the model was linearised about the robot’s initial position and commanded to follow a minimum-jerk trajectory. In this case too, the robot was unable to track the trajectory and eventually fell.
IV-D3 Multi-rate study
@@ V-A Linearization
Exploiting the property S(x)y=S(y)xS(x)y=-S(y)x and writing $^{B}\omega_{B,J_{j}}=J^{rel}_{\omega_{j}}(s)\dot{s}$ we can finally rewrite:
@@
where $^{B}r_{i}$ is the distance between the centre of mass and the $i^{th}$ turbine expressed in body frame. Exploiting once again the definition of the derivative of the rotation matrix, the property S(x)y=S(y)xS(x)y=-S(y)x and using the linear and angular Jacobians $^{B}\omega_{B,J_{j}}=J^{rel}_{\omega_{j}}(s)\dot{s}$ and $^{B}\omega_{B,J_{j}}=J^{rel}_{\omega_{j}}(s)\dot{s}$, (27) rewrites as: